SpaceRef

SpaceRef


NASA Cassini Significant Events 03/09/11 - 03/15/11

Status Report From: Jet Propulsion Laboratory
Posted: Saturday, March 19, 2011

image The most recent spacecraft telemetry was acquired on Mar. 15 from the Deep Space Network tracking complex at Goldstone, California. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Information on the present position and speed of the Cassini spacecraft may be found on the "Present Position" page at: http://saturn.jpl.nasa.gov/mission/presentposition/.

Wednesday, March 9 (DOY 068)

An early look at the Reaction Wheel Assembly (RWA) performance so far in S67 indicated that RWA-3 continues to be nominal and is running even smoother than expected. RWA-3 has been on-line for eight days now. The wheel performance will continue to be closely monitored throughout the S67 sequence.

Thursday, March 10 (DOY 069)

The final sequence activity performed for S66 this week was the release of the S66 as-flown sequence products. As-flown products are generated following the completion of execution of each sequence. S66 completed execution on Mar. 7.

Friday, March 11 (DOY 070)

Real-time command files were uplinked to the spacecraft in support of S67 for the clearing of the Command and Data Subsystem (CDS) uplink log, and for the Magnetometer (MAG) power cycle to recover from a Bus Interface Unit (BIU) sick state on DOY-070 at 05:53:00. MAG has recovered and is in its normal operating state.

As part of the ongoing task of transitioning Cassini off of the Andrew File System (AFS) to the Network File System (NFS), user and application testing of the Distributed Object Manager (DOM) on NFS began this week. DOM NFS testing resulted in a handful of issues. Further investigation to resolve issues uncovered during the testing is ongoing.

Saturday, March 12 (DOY 071)

The Cosmic Dust Analyzer (CDA) team performed a decontamination of the instrument's detector in S67 which executed normally. The instrument reset and flight software re-load from the Solid State Recorder (SSR) went well; the decontamination reached thermal equilibrium. Decontaminations occur every few months and are part of normal CDA instrument maintenance. Based on current telemetry, the instrument status is normal.

Sunday, March 13 (DOY 072)

This week in science, as the spacecraft was on its inbound leg toward the Rev 146 Saturn periapsis, the observations continued from last week with a focus on Saturn. The Optical Remote Sensing (ORS) instruments performed a sequence of observations to study the atmosphere of Saturn. The Imaging Science Subsystem (ISS) and the Visual and Infrared Mapping Spectrometer (VIMS) took turns making observations to learn more about Saturn's wind speeds and the giant storm which was first observed in early December 2010, and is continuing. The Composite Infrared Spectrometer (CIRS) measured abundances and distribution of oxygen and molecular compounds in Saturn's stratosphere as a function of latitude. The Ultraviolet Imaging Spectrograph (UVIS) made an Extreme Ultraviolet/Far Ultraviolet map of the planet to study the upper atmosphere. Lastly, CIRS took a Saturn mid-infrared map for the purpose of determining Saturn's upper troposphere and tropopause temperatures, and the Navigation Team took some images of Hyperion for optical navigation. The ORS teams also performed two Titan cloud monitoring observations.

Monday, March 14 (DOY 073)

A teleconference with the science team leads, typically held monthly during the Solstice Mission, was held today to discuss the Decadal Survey 2013 - 2022 results which were recently released, and to recap the events of the 42nd Lunar and Planetary Science Conference (LPSC) attended by Cassini scientists and Outreach team members. The LPSC, jointly sponsored by the Lunar and Planetary Institute and NASA Johnson Space Center, brings together international specialists in petrology, geochemistry, geophysics, geology and astronomy to present the latest results of research in planetary science. Since its beginning in 1970, the LPSC has been a significant focal point for planetary science research, with more than 1500 planetary scientists and students attending from all over the world. The 42nd LPSC was held March 7-11, 2011, in The Woodlands, Texas.

Port 2 products were due today as part of the S69 Sequence Implementation Process (SIP). The products will be merged and sent out to the flight team for review.

Tuesday, March 15 (DOY 074)

Mission Planning held a "Proximal Orbits Planning Kickoff" meeting today. Spacecraft Operations, Navigation, and Science Planning and Sequencing Team members supported the meeting in an effort to start looking at the proximal orbits and determine what work needs to be done, and what resources will be needed to execute this phase of the mission.

The command loss timer (CLT) value was changed temporarily to 125 hours today, and will be restored back to its nominal default value of 115 hours on Mar. 19.


// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.